OpenStudy (anonymous):

Find the first six terms of the sequence. a1 = -1, an = 2 • an-1

OpenStudy (math&ing001):

Hey there @lornbeach , All you have to do is plug in the values of an-1 to find the next term. Lets start with a2 = 2 . a1 = 2 * (-1) = -2 ; a3= 2 . a2 = 2 * (-2) = -4 and so on. Tell me if you need help with the rest.